OpenStudy (anonymous):

multiply and simplify exponential notation with positive exponents (6v)5 * (6v)7

OpenStudy (anonymous):

multiply and simplify exponential notation with positive exponents (6v)5 * (6v)7 is the anse 1260v2

OpenStudy (anonymous):

Am I right in assuming that the equation should look like\[(6v)^5*(6v)^7\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you've got exponents, you simply add them. I'll prove to you why this is the case with your question: \[(6v)^5=6v*6v*6v*6v*6v\]\[(6v)^7=6v*6v*6v*6v*6v*6v*6v\]So if you multiply these together, you simply get 6v multiplied by itself 12 times, or (6v)^12

OpenStudy (anonymous):

Because 5+7, and if you define exponents to be a notation showing how much something is MULTIPLIED, then if you multiply more of the same to a particular term you simply have to add the exponents. I.e.,\[a^1*a^1=a^2=a*a\]

OpenStudy (anonymous):

Sorry, I meant to say 5+7=12

OpenStudy (anonymous):

Also, if you want to denote an exponent you use a ^ sign, followed by the exponent itself (SHIFT + 6 on your keyboard, probably).

OpenStudy (anonymous):

tangent so the answer would be 12

OpenStudy (anonymous):

(6v)^12. I can't write it with the equation function on this forum because it only lets me put 1 item in the exponent place, but yes, 6v*6v*6v*6v*6v*6v*6v*6v*6v*6v*6v*6v, or (6v)^12

OpenStudy (anonymous):

so i can write the answer (6v)^12 keeping the 6v in parenthese

OpenStudy (anonymous):

Yes. If you wrote 6v^12, that would mean 6*(v)^12, certainly not what you had at the beginning.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem, Jazi. Stop by another time.

OpenStudy (anonymous):

divide and wite using nonnegative exponents the quotient is--? 30x^4y^2 / 5x^7y^2 =

OpenStudy (anonymous):

checking my answer plse divide and wite using nonnegative exponents the quotient is--? 30x^4y^2 / 5x^7y^2 = the quotient is 6/x^3 is this correct